Identify the slope and y-intercept of the line graph the equation 3x+Y=-1
Answer by jim_thompson5910(35256)

Solve for y.


3x+y=-1


y=-1-3x


y=-3x-1


The equation y=-3x-1 is in y=mx+b form where the slope is m=-3 and the y-intercept is b=-1. This is the point (0,-1)